# ME-Switch: A Memory-Efficient Expert Switching Framework for Large Language Models

### v1

@@ Abstract
The typical process for LLM’s development involves pre-training a general foundation model on massive data, followed by fine-tuning on task-specific data to obtain a series of specialized experts. Serving these experts can pose significant challenges, as loading all experts onto devices is impractical, and frequent switching between experts in response to user requests can incur substantial I/O costs, leading to a notable increase in latency and serving expenses. Previous approaches decompose the expert weights into those of the pre-trained model and residual delta weights, followed by quantizing the delta weights using output channel-wise step sizes to reduce the model size. However, these methods overlook the fact that certain input channels of delta weights can cause significant quantization errors at extremely low bitwidths. Additionally, existing methods assume that the appropriate model for a user request is known in advance, which is not the case in practice. To this end, we introduce ME-Switch, a memory-efficient expert switching framework tailored for LLM serving. To condense the number of bits required for describing the delta weights, ME-Switch employs mixed-precision quantization that selectively quantizes non-salient input channels of delta weights to extremely low bits while keeping the salient ones intact, which significantly reduces storage demand while maintaining performance. Moreover, we develop a routing method that efficiently directs user queries to the most suitable expert by transforming the model selection problem into a domain classification problem. Extensive experiments show the promising memory efficiency and routing performance of ME-Switch. For example, when serving three models from Mistral-7B family, ME-Switch reduces the model size by 1.74×1.74$\times$1.74 × and maintains nearly lossless performance on instruction, mathematical reasoning and code generation tasks. Furthermore, our method can efficiently serve 16 models from the Mistral-7B family on a NVIDIA A100 GPU.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 [6] and Gemini [43], have achieved significant advancements in natural language processing (NLP). Trained on large-scale text datasets, these models develop a broad foundation of language understanding, enabling them to excel at tasks requiring common-sense knowledge. To acquire task-specific knowledge, LLMs can be further fine-tuned on specialized tasks, thereby adapting them for diverse applications such as interactive agents [45, 22], code generation [34, 31], and mathematical problem solving [33, 31], showing the remarkable versatility of LLMs. Nevertheless, given that no single model can master all tasks simultaneously, serving multiple LLMs, each precisely tailored for specific tasks, becomes essential for good performance.
@@
To address the above challenges, existing methods [28, 50] decompose the weights of fine-tuned models into the weights of the pre-trained model and the delta weights introduced during fine-tuning. By applying per-tensor quantization [35] to these delta weights, it is possible to significantly reduce storage requirements, thereby enabling efficient sharing of the base model’s storage across multiple fine-tuned models. Nevertheless, using per-tensor quantization ignores the distinction in delta weight values across different input and output channels, resulting in substantial information loss. To mitigate this issue, finer granularity quantization, such as output channel-wise quantization [48, 47, 27] where each output channel is allocated its own learnable step size, still neglects the variations in input channels, as shown in Figure 1. To further mitigate information loss, one may rescale the input channels of delta weights before quantization [26]. However, with extremely low-bitwidth quantization, rescaling provides only limited alleviation. Moreover, existing methods assume that the appropriate model to utilize in response to a user request is known in advance. In reality, determining which model to use dynamically is crucial for handling diverse user queries efficiently. This gap highlights the need for an intelligent model selection mechanism that can adaptively choose the best-suited model based on the specific requirements of each user request.
@@
In this paper, we propose ME-Switch, a memory-efficient expert switching framework tailored for LLMs. To reduce the quantization error while simultaneously reducing storage needs, we develop a mixed-precision quantization method that quantizes non-salient input channels of delta weights to extremely low bits while preserving those salient ones, which could substantially increase quantization errors when quantized at very low bitwidths, in full precision. Since the number of salient input channels is relatively small, incorporating a limited amount of high-precision delta weights incurs negligible memory overhead and inference cost. To identify the important input channels of delta weights, one may select based on their magnitudes [13], as shown in Figure 2(a), which however fails to capture those leading to high quantization error. In contrast, our approach identifies salient input channels of delta weights based on their impact on reconstruction errors in the output activations, as illustrated in Figure 2(b). Additionally, we develop a routing method that selects the most suitable LLM in response to a user request, enabling efficient and accurate model switching. To simplify the routing problem, we assume that each LLM specializes in a distinct domain, such as code generation or mathematics reasoning. This allows us to treat model selection as a multiple-choice question-answering task where each option represents a target domain, effectively transforming the problem into a domain classification task, as shown in Figure 3. Notably, dialogue LLMs exhibit strong instruction-following capabilities. Leveraging this, we propose using a small pre-trained dialogue LLM as a model-level router to solve the multiple-choice question-answering task. We then construct a multiple-choice question-answering dataset and fine-tune the router to improve its performance. Given its smaller size compared to the specialized LLMs, the router operates efficiently without adding significant overhead.
@@
Our contributions can be summarized as follows. 1) We propose a mixed-precision quantization method that significantly reduces the storage demands of serving multiple LLMs while maintaining performance. This is achieved by selectively quantizing the non-salient input channels of delta weights, leaving the salient ones unchanged. 2) We introduce a routing method that dynamically selects the most appropriate LLM for a given query, ensuring efficient and accurate model switching. This is accomplished by transforming the model selection problem into a domain classification task and then solving it with a small LLM. 3) We conduct extensive experiments demonstrating the promising memory efficiency and routing performance of ME-Switch. Remarkably, when serving three models from the Mistral-7B family, ME-Switch not only delivers near-lossless performance on instruction, mathematical reasoning, and code generation tasks but also reduces the model size by 1.74×1.74$\times$1.74 ×. Even more impressively, our method is able to simultaneously serve up to $16$ models for Mistral-7B family on a single NVIDIA A100 GPU without encountering the out-of-memory issue.
@@ 2 Related Work
Efficient LLM serving. LLMs can be efficiently deployed on GPUs for high-throughput serving using several inference frameworks, such as vLLM [24] and Orca [51]. Given that a single LLM cannot excel across all domains, it is crucial to serve multiple LLMs simultaneously to handle diverse user queries effectively. To determine which model to use during inference, Zooter [32] utilizes a task-level router to load different LLMs based on the incoming task requirements. This approach, however, introduces significant memory challenges, as hosting all LLMs on a GPU simultaneously can excessively strain the available VRAM. One solution is to serve multiple Low-Rank Adaptation (LoRA) modules within a multi-tenant serving system, such as Punica [9] and S-LoRA [38]. Another viable strategy to address the above issue involves dynamically loading different LLMs from CPU memory to GPU memory as needed, thereby reducing peak GPU memory utilization. Nevertheless, this dynamic loading competes for GPU memory bandwidth with model-swapping operations, presenting a significant bottleneck. To mitigate these challenges, techniques like Deltazip [50] and BitDelta [28] have been developed to compress the delta parameters. These methods allow task-specific delta parameters to be loaded into GPU memory on-demand, ensuring that only a single pre-trained model’s parameters reside permanently in GPU memory. This approach aims to achieve low-latency inference while minimizing the costs associated with maintaining numerous fine-tuned models in GPU memory. Our approach stands out by serving multiple LLMs with a minimal memory footprint with nearly lossless performance, which is achieved through an advanced delta weights quantization strategy and an adaptive model-level router.
@@
Delta compression. In recent years, numerous approaches have focused on reducing the storage overhead for maintaining different task-specific models through delta parameter compression. Model merging strategies often incorporate multiple tasks’ delta parameters [14] into the pretrained parameters to minimize the number of parameters needed for multi-task operations. To address the parameter conflicts that often arise during model merging, Yu et al. [53] proposed a method involving massive unstructured random pruning (achieving 90% sparsity) of delta parameters and subsequent rescaling. This approach ensures that the accuracy of downstream tasks is not compromised. In a similar vein, Ties-merging [49] developed a pruning strategy based on the magnitude and sign of delta parameters to further reduce conflicts and storage needs. Additionally, the LoRA-based PEFT methods [19, 46] introduces a novel approach by learning one or several low-rank matrices to represent the delta parameters. ZipLoRA [37] explores the sparsity within these low-rank matrices, allowing the fusion of low-rank matrices from different tasks to further reduce the spatial overhead of multi-task models. However, these merged models often suffer from notable performance degradation. Conversely, some studies focus solely on compressing delta parameters for each task without merging, thereby mitigating performance degradation across multiple tasks. For example, Ryu et al. [36] combine quantization and low-rank estimation techniques to reduce the storage size of delta parameters. Liu et al. [28] push this further by quantizing each task’s delta parameters to 1 bit, achieving more than a tenfold compression. However, these quantization methods often lack robust handling of outliers, resulting in performance declines compared to uncompressed delta parameters. Moreover, models that do not integrate multi-task delta parameters require manual activation of specific delta parameters for each task, which reduces the model’s applicability in multi-task environments.
@@ 3 Preliminaries
For simplicity, we employ uniform quantization [21] to compress the models. Given a matrix ${\bf X}$ with floating-point values (e.g., FP16 or FP32), the quantization process can be expressed as:
@@
Recent studies [28, 50] have shown that the weights of a fine-tuned model can be decomposed into the weights of the pre-trained model and the delta weights introduced during fine-tuning. Let ${\bf W}\in\mathbb{R}^{m\times n}$ and ${\bf W}_{\mathrm{FT}}\in\mathbb{R}^{m\times n}$ be the weight matrices of the pre-trained model and the fine-tuned model, respectively, where $m$ represents the number of input channels and $n$ denotes the output number of channels. The delta weight is defined as $\Delta={\bf W}_{\mathrm{FT}}-{\bf W}$. To reduce storage requirements, one can perform quantization using Eq. (1) to compress $\Delta$. However, this method often results in significant performance degradation because it assumes all delta weight channels are equally sensitive to quantization noise. Additionally, existing methods assume the appropriate model for a user request is predetermined. In practice, dynamically choosing the most suitable model based on the specific nature of each user’s query is more effective.
@@ 4 Proposed Method
In this section, we propose ME-Switch, a memory-efficient expert switching framework designed to optimize the deployment of LLMs. ME-Switch addresses the challenges of reducing storage demands while maintaining performance and introduces model-level routing to efficiently manage task-specific queries. We will introduce the details of ME-Switch in the following.
@@ 4.1 Salient-Aware Delta Compression
In this section, we introduce our salient-aware delta compression approach, which specifically targets the preservation of salient input channels during quantization. While quantization methods with learnable step sizes for each output channel can handle variations in output channels effectively [48, 27], variations in input channels pose a significant challenge to maintaining model performance after quantization, as shown in Figure 1. Some salient input channels of delta weights, regardless of their magnitude, can cause substantial information loss and degrade model performance when quantized to a very low bitwidth. To address this, rescaling the input channels of delta weights before quantization [26] offers a solution to mitigate some of the quantization error, which however provides only limited alleviation under the context of extremely low-bit quantization.
Salient-aware delta selection. To protect salient input channels of delta weights, we develop a mixed-precision quantization method, where the majority of input channels of delta weights are quantized to low-bit precision, while only a small number of critical input channels are represented in 16-bit. This approach achieves significant reductions in storage requirements with minimal performance loss. The remaining challenge lies in identifying these salient input channels. A naive approach might involve selecting the input channels with large magnitudes as the salient delta weights [13], as shown in Figure 2(a). However, this method neglects the influence of input activations on the outputs, failing to identify channels that lead to high quantization error. To address this, inspired by the pruning metric [41], we introduce a new salient delta weights selection metric based on reconstruction errors in the outputs, considering both weights and input activations, as shown in Figure 2(b). Specifically, given an input ${\bf x}\in\mathbb{R}^{m}$, the reconstruction error for the input channel $i$ can be calculated by
@@
Efficient distillation. The quantization step size defined in Eq. (1) plays an important role in the final performance. To learn the quantization step size, we employ knowledge distillation to guide the alignment of the output logits of the quantized model with those of the full-precision fine-tuned model following [28]. To reduce training overhead, we freeze the model weights and focus solely on optimizing the quantization step size ${\bf s}$ using a small calibration dataset $\mathcal{X}$ by solving problem
@@
On-demand swapping. The extremely low-bit compression for delta weights significantly reduces the model size, thereby alleviating storage demands. This approach enables us to maintain a single pre-trained model while storing multiple sets of compressed delta weights, facilitating efficient on-demand swapping. In this scenario, the pre-trained model remains in GPU memory, and the corresponding compressed delta weights are loaded dynamically based on the user query. Compared with directly serving multiple models, our method is more GPU memory-efficient. Since LLM decoding is memory-bound [29, 39] due to the auto-regressive nature, reducing model size effectively decreases the parameter loading time, thereby improving decoding latency. To achieve fast inference, we decouple the matrix multiplication during inference into two components:
@@
where $\tilde{\Delta}$ represents the compressed delta weights, including both the quantized unsalient and the full-precision salient delta weights. For ${\bf x}{\bf W}$, the computation is performed using a FP16 batched GEMM kernel. For ${\bf x}\tilde{\Delta}$, we implement an efficient Triton kernel [44] that fuses dequantization and matrix multiplication for efficient computation.

### v2

@@ Abstract
The typical process for LLM’s development involves pre-training a general foundation model on massive data, followed by fine-tuning on task-specific data to obtain a series of specialized experts. Serving these experts can pose significant memory challenges, as loading all experts onto devices is impractical, and frequent switching between experts in response to user requests can incur substantial I/O costs. Previous approaches decompose the expert weights as the pre-trained weights plus delta weights, followed by quantizing the delta weights using output channel-wise step sizes to reduce the model size. However, these methods overlook the fact that certain input channels of delta weights can cause significant quantization errors at extremely low bitwidths. Additionally, existing methods assume that the appropriate model for a user request is known in advance, which is not the case in practice. To this end, we introduce ME-Switch, a memory-efficient expert switching framework tailored for serving multiple LLMs. To condense the number of bits required for describing the delta weights, we propose a salient-aware delta compression method that first identifies which input channels of delta weights are salient based on reconstruction error and then employs mixed-precision quantization that selectively quantizes non-salient input channels of delta weights to extremely low bits while keeping the salient ones intact, significantly reducing storage demand while maintaining performance. Moreover, we develop a model-level routing method that efficiently directs user queries to the most suitable expert by performing domain classification. Extensive experiments show the promising memory efficiency and routing performance of ME-Switch. For example, when serving three models from the Mistral-7B family, ME-Switch reduces the model size by 1.74×1.74$\times$1.74 × and maintains nearly lossless performance on instruction, mathematical reasoning, and code generation tasks. Furthermore, our method can efficiently serve 16 Mistral-7B models on an NVIDIA A100 GPU.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), have achieved significant advancements in natural language processing (NLP). Trained on large-scale text datasets, these models develop a broad foundation of language understanding, enabling them to excel at tasks requiring common-sense knowledge. To acquire task-specific knowledge, LLMs can be further fine-tuned on specialized tasks, thereby adapting them for diverse applications such as interactive agents (Touvron et al., 2023; Jiang et al., 2023), code generation (Luo et al., 2023b; Lozhkov et al., 2024), and mathematical problem solving (Luo et al., 2023a; Lozhkov et al., 2024), showing the remarkable versatility of LLMs. Nevertheless, given that no single model can master all tasks simultaneously, serving multiple LLMs, each precisely tailored for specific tasks, becomes essential for good performance.
@@
To address the above challenges, existing methods (Liu et al., 2024a; Yao & Klimovic, 2023) decompose the weights of fine-tuned models into pre-trained weights and delta weights introduced during fine-tuning. While low-rank approximation (Hu et al., 2022) can compress these delta weights, it falls short for full fine-tuned models whose delta weights lack low-rank properties (Liu et al., 2024a; Lialin et al., 2024; Hao et al., 2024). As a result, prior work has adopted per-tensor quantization (Liu et al., 2024a) as an alternative, significantly reducing storage needs and facilitating efficient sharing of the base model’s storage across multiple models. Nevertheless, this method ignores the distinction in delta weight values across different input and output channels, resulting in substantial information loss. To mitigate this issue, output channel-wise quantization (Xiao et al., 2023; Wei et al., 2023; Liu et al., 2024c) where each output channel is allocated its own learnable step size, still neglects input channel variations, as shown in Figure 1. To further mitigate information loss, rescaling input channels before quantization can be employed, but this provides only limited alleviation at extremely low bitwidths. In addition, existing methods assume that the appropriate model to utilize in response to a user request is known in advance. In reality, dynamically selecting the optimal model is essential for effectively handling diverse queries. This gap highlights the need for an intelligent model selection mechanism that can identify the most suitable model based on specific user needs.
@@
In this paper, we propose ME-Switch, a memory-efficient expert switching framework tailored for LLMs. To reduce the quantization error while simultaneously reducing storage needs, we develop a mixed-precision quantization method that quantizes non-salient input channels of delta weights to extremely low bits while preserving those salient ones, which could substantially increase quantization errors when quantized at very low bitwidths, in full precision. Since the number of salient input channels is relatively small, incorporating a limited amount of high-precision delta weights incurs negligible memory overhead and inference cost. To identify the important input channels of delta weights, one may select based on their magnitudes (Dettmers et al., 2022), as shown in Figure 2(a), which however fails to capture those leading to high quantization error. In contrast, our approach identifies salient input channels of delta weights based on their impact on reconstruction errors in the output activations, as illustrated in Figure 2(b). Additionally, we develop a routing method that selects the most suitable LLM in response to a user request, enabling efficient and accurate model switching. To simplify the routing problem, we assume that each LLM specializes in a distinct domain, such as code generation or mathematics reasoning. This allows us to treat model selection as a multiple-choice question-answering task where each option represents a target domain, effectively transforming the problem into a domain classification task, as shown in Figure 3. Notably, dialogue LLMs exhibit strong instruction-following capabilities. Leveraging this, we propose using a small pre-trained dialogue LLM as a model-level router to solve the multiple-choice question-answering task. We then construct a multiple-choice question-answering dataset and fine-tune the router to improve its performance. Given its smaller size compared to the specialized LLMs, the router operates efficiently without adding significant overhead.
@@
Our contributions can be summarized as follows. 1) We introduce ME-Switch, a memory-efficient framework designed for serving multiple LLMs. It only stores a single full-precision pre-trained model and dynamically loads the appropriate compressed delta model weights in response to user queries. 2) We develop a mixed-precision quantization method that significantly reduces the storage demands of serving multiple LLMs while maintaining performance, which is achieved by selectively quantizing the non-salient input channels of delta weights and leaving the salient ones unchanged. We further introduce a routing method that dynamically selects the most appropriate LLM for a given query, which is accomplished by transforming the model selection problem into a domain classification task and then solving it with a small LLM. 3) We conduct extensive experiments demonstrating the promising memory efficiency and routing performance of ME-Switch. Remarkably, when serving three models from the Mistral-7B family, ME-Switch not only delivers near-lossless performance on instruction, mathematical reasoning, and code generation tasks but also reduces the model size by 1.74×1.74$\times$1.74 ×. More impressively, our method is able to serves up to 16 Mistral-7B models on a single NVIDIA A100 GPU without running out of memory.
@@ 2 Related Work
Efficient LLM serving. LLMs can be efficiently deployed on GPUs for high-throughput serving using several inference frameworks, such as vLLM (Kwon et al., 2023) and Orca (Yu et al., 2022). Given that a single LLM cannot excel across all domains, it is crucial to serve multiple LLMs simultaneously to handle diverse user queries effectively. To determine which model to use during inference, Zooter (Lu et al., 2023) utilizes a task-level router to load different LLMs based on the incoming task requirements. This approach, however, introduces significant memory challenges, as hosting all LLMs on a GPU simultaneously can excessively strain the available VRAM. One solution is to serve multiple Low-Rank Adaptation (LoRA) modules within a multi-tenant serving system, such as Punica (Chen et al., 2023) and S-LoRA (Sheng et al., 2023a). However, these methods fail to accommodate full fine-tuned models because such models typically do not display the low-rank characteristics necessary for LoRA approximation. Another viable strategy to address the above issue involves dynamically loading different LLMs from CPU memory to GPU memory as needed, thereby reducing peak GPU memory utilization. Nevertheless, this dynamic loading competes for GPU memory bandwidth with model-swapping operations, presenting a significant bottleneck. To mitigate these challenges, techniques like Deltazip (Yao & Klimovic, 2023) and BitDelta (Liu et al., 2024a) have been developed to compress the delta parameters. These methods allow task-specific delta parameters to be loaded into GPU memory on-demand, ensuring that only a single pre-trained model’s parameters reside permanently in GPU memory. This approach aims to achieve low-latency inference while minimizing the costs associated with maintaining numerous fine-tuned models in GPU memory. Our approach stands out by serving multiple LLMs with a minimal memory footprint with nearly lossless performance, which is achieved through an advanced delta weights quantization strategy and an adaptive model-level router.
@@
Delta compression. In recent years, numerous approaches have focused on reducing the storage overhead for maintaining different task-specific models through delta parameter compression. Model merging strategies often incorporate multiple tasks’ delta parameters (Ding et al., 2023) into the pretrained parameters to minimize the number of parameters needed for multi-task operations. To address the parameter conflicts that often arise during model merging, Yu et al. (Yu et al., 2023a) proposed a method involving massive unstructured random pruning (achieving 90% sparsity) of delta parameters and subsequent rescaling. This approach ensures that the accuracy of downstream tasks is not compromised. In a similar vein, Ties-merging (Yadav et al., 2024) developed a pruning strategy based on the magnitude and sign of delta parameters to further reduce conflicts and storage needs. Additionally, the LoRA-based PEFT methods (Hu et al., 2021; Valipour et al., 2022; Ping et al., 2024) introduce a novel approach by learning one or several low-rank matrices to represent the delta parameters. ZipLoRA (Shah et al., 2023) explores the sparsity within these low-rank matrices, allowing the fusion of low-rank matrices from different tasks to further reduce the spatial overhead of multi-task models. However, these methods struggle to accurately approximate delta parameters in full fine-tuned models, as these models usually lack low-rank properties (Liu et al., 2024a; Lialin et al., 2024; Hao et al., 2024). Conversely, some studies focus solely on compressing delta parameters for each task without merging, thereby mitigating performance degradation across multiple tasks. For example, Ryu et al. (Ryu et al., 2023) combine quantization and low-rank estimation techniques to reduce the storage size of delta parameters. Liu et al. (Liu et al., 2024a) push this further by quantizing each task’s delta parameters to 1 bit, achieving more than a tenfold compression. However, these quantization methods often lack robust handling of outliers, resulting in performance declines compared to uncompressed delta parameters. Moreover, models that do not integrate multi-task delta parameters require manual activation of specific delta parameters for each task, which reduces the model’s applicability in multi-task environments.
@@ 3 Preliminaries
For simplicity, we employ uniform quantization (Jacob et al., 2018) to compress the models. Given a matrix ${\bf X}$ with floating-point values (e.g., FP16), the quantization process can be expressed as:
@@
Recent studies (Liu et al., 2024a; Yao & Klimovic, 2023) have shown that the weights of a fine-tuned model can be decomposed into the weights of the pre-trained model and the delta weights introduced during fine-tuning. Let ${\bf W}\in\mathbb{R}^{m\times n}$ and ${\bf W}_{\mathrm{FT}}\in\mathbb{R}^{m\times n}$ be the weight matrices of the pre-trained model and the fine-tuned model, respectively, where $m$ represents the number of input channels and $n$ denotes the output number of channels. The delta weight is defined as $\Delta={\bf W}_{\mathrm{FT}}-{\bf W}$. To reduce storage requirements, one can perform quantization using Eq. (1) to compress $\Delta$. However, this method often results in significant performance degradation because it assumes all delta weight channels are equally sensitive to quantization noise. Additionally, existing methods assume the appropriate model for a user request is predetermined. In practice, dynamically choosing the most suitable model based on the specific nature of each user’s query is more effective.
@@ 4 Proposed Method
In this section, we propose ME-Switch, a memory-efficient expert switching framework designed to optimize the deployment of multiple LLMs. ME-Switch reduces memory demands while maintaining performance by using mixed-precision quantization, which quantizes non-salient weights to extremely low bitwidths while keeping salient weights intact. Additionally, ME-Switch introduces model-level routing to efficiently handle task-specific queries. We will detail ME-Switch in the following sections.
@@ 4.1 Salient-Aware Delta Compression
In this section, we introduce our salient-aware delta compression approach, which specifically targets the preservation of salient input channels during quantization. While quantization methods with learnable step sizes for each output channel can handle variations in output channels effectively (Xiao et al., 2023; Liu et al., 2024c), variations in input channels pose a significant challenge to maintaining model performance, as shown in Figure 1. Some salient input channels of delta weights, regardless of their magnitude, can cause substantial information loss and degrade model performance when quantized to a very low bitwidth. To address this, rescaling the input channels of delta weights before quantization (Lin et al., 2024) offers a solution to mitigate some of the quantization error, which however provides only limited alleviation under the context of extremely low-bit quantization.
Salient-aware delta selection. To protect salient input channels of delta weights, we develop a mixed-precision quantization method, where the majority of input channels of delta weights are quantized to low-bit precision, while only a small number of critical input channels are represented in their original precision. This approach achieves significant reductions in storage requirements with minimal performance loss. The remaining challenge lies in identifying these salient input channels. A naive approach might involve selecting the input channels with large magnitudes as the salient delta weights (Dettmers et al., 2022), as shown in Figure 2(a). However, this method neglects the influence of input activations on the outputs, failing to identify channels that lead to high quantization error. To address this, inspired by the pruning metric (Sun et al., 2024), we introduce a salient delta weights selection metric based on reconstruction errors in the outputs, considering both weights and input activations, as shown in Figure 2(b). Specifically, given an input ${\bf x}\in\mathbb{R}^{m}$, the reconstruction error for the input channel $i$ can be calculated by
@@
Efficient distillation. The quantization step size in Eq. (1) plays an important role in the final performance. To learn the quantization step size, we employ knowledge distillation to guide the alignment of the output logits of the quantized model with those of the full-precision fine-tuned model following (Liu et al., 2024a). To reduce training overhead, we freeze the delta weights and focus solely on optimizing the quantization step size ${\bf s}$ using a small calibration dataset $\mathcal{X}$ by solving
@@
On-demand swapping. The extremely low-bit compression for delta weights significantly reduces the model size, alleviating storage demands. This approach enables us to maintain a single pre-trained model while storing multiple sets of compressed delta weights, facilitating efficient on-demand swapping. In this scenario, the pre-trained model remains in GPU memory, and the corresponding compressed delta weights are loaded dynamically based on the user query. Compared with directly serving multiple models, our method is more GPU memory-efficient. Since LLM decoding is memory-bound (Liu et al., 2023; Sheng et al., 2023b) due to the auto-regressive nature, reducing model size effectively decreases the parameter loading time, thereby improving decoding latency. To achieve fast inference, we decouple the matrix multiplication during inference into two components:
@@
where $\tilde{\Delta}$ represents the compressed delta weights, including both the quantized unsalient and the full-precision salient delta weights. For ${\bf x}{\bf W}$, the computation is performed using a FP16 batched GEMM kernel. For ${\bf x}\tilde{\Delta}$, we implement an efficient Triton kernel (Tillet et al., 2019) that fuses dequantization and matrix multiplication for efficient computation. The reduced memory footprint of the compressed delta weights enables our method to perform batched forward passes across multiple models simultaneously. This batching at the model level significantly enhances efficiency compared to the traditional approach, which processes each model individually—especially beneficial when serving multiple models (See Figure 8).
4.2 Model-level Routing
Given a user query, we present a method to determine the appropriate model. Consider a set of LLMs represented as =$\mathcal{F}=\{f_{1},f_{2},\cdots,f_{M}\}$, where $M$ denotes the number of models. Given a user query $q$, we aim to find the most suitable LLM by solving the following problem $\arg\max_{f\in\mathcal{F}}P(q,f(q))$, where $P$ is a function that measures the quality or performance of the LLM response. To simplify the routing process, we assume that each LLM in $\mathcal{F}$ specializes in distinct domains such as code generation or mathematical problem solving. This setup allows us to treat the routing challenge as a multiple-choice question-answering task, where each option corresponds to a specific domain, thereby transforming the problem into a domain classification problem. Note that dialogue LLMs like Qwen1.5-1.8B-Chat (Bai et al., 2023) exhibit capabilities in following instructions, which inspires us to utilize a small pre-trained LLM as a model-level router. As illustrated in Figure 3, we first prompt the router with the user’s question using a template designed to elicit domain classification. For the prompt template, please refer to Section A of the appendix. Based on the router’s response, we then dynamically load the corresponding compressed delta weights for the selected domain-specific model, such as a mathematical model, to generate outputs.